OpenStudy (anonymous):

What are the first 5 terms of the sequence an = n − 6? please help D;

OpenStudy (anonymous):

an = n -6 put n=1 for first term a1 = 1-6 = -5 similarly put n= 2,3,4,5 for the other 4 terms

OpenStudy (anonymous):

so it just -6, -5, -4, -3, -2,?

OpenStudy (anonymous):

the first 5 terms of this sequence can be found by plugging in n= 1,2,3,4,5 ( because you are only finding the first 5 terms) an=n-6 a1= 1- 6 a2=2-6 a3=3-6 Your 5 terms are now : 5 , 4 , 3, 2 , 1 a4=4-6 a5=5-6